OpenStudy (anonymous):

Need help with solving systems of equations. The equations are: 0.3x-0.2y=4 0.4x+0.5y=63/19

OpenStudy (anonymous):

We can cancel out either x or y

OpenStudy (anonymous):

yep, either or.

OpenStudy (anonymous):

Let's cancel out y To do so we multiply first equation by 5 and second equation by 2

OpenStudy (anonymous):

need to cancel out those icky little decimals first it seems. I'm just getting horribly wrong answers by myslef

OpenStudy (anonymous):

We get 1.5x-y=20 .8x+y=126/19

OpenStudy (anonymous):

Getting rid of an icky decimal is as easy as multiplying every term by 10.

OpenStudy (anonymous):

When we add two equation, y cancel; leaving us with 2.3x=506/19

OpenStudy (anonymous):

yep mooby, it seems I had the steps in order but forgot to multiply my fraction as well. trying it out myself now to see if I get what you guys get

OpenStudy (anonymous):

@Mooby true but why work with big numbers if you can work with little ones?

OpenStudy (anonymous):

I like whole numbers. ;)

OpenStudy (anonymous):

2.3 = 23/10 :-)

OpenStudy (anonymous):

hmm getting x = 220/19

OpenStudy (anonymous):

so thats the same as imranmeah91's 2.3x=506/19 yea?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

yay! now I can prey I get the y right. Im absolutely horrible at the elimination method, and substitution wasn't looking any better.

OpenStudy (anonymous):

if you divide both side by 2.3, you would get same answer

OpenStudy (anonymous):

Im at -50/19 = y sound right?

OpenStudy (anonymous):

0.3x-0.2y=4 0.4x+0.5y=63/19 y=4-0.3x 0.4x+0.5(4-0.3x)=63/19 0.4x+2-0.15x=63/19 0.4x+0.15x=(63/19)-2 0.55x=(63/19)-2 x=((63/19)-2)/0.55 y=4-0.3(((63/19)-2)/0.55)

OpenStudy (anonymous):

yep its right :) got a correct on ym school webbie. Thanks for the help guys/gals!

OpenStudy (anonymous):

I would multiply first equation by 10 3x-2y=40 3(220/19)-2y=40 y=-50/19